A, B & C form the vertices of a triangle.

 CAB = 90°, ∠

 ABC = 74° and AB = 8.9.

Calculate the length of BC rounded to 3 SF.


1
Expert's answer
2021-05-24T00:24:01-0400

In triangle BAC

"\angle A=90^o\\
\angle B=74^o\\"

And AB= 8.9

So

"Cos(74^o)=\frac{AB}{BC}"

"0.27=\frac{8.9}{BC}"


BC=32.9